5. (12pts) Find the exact value of the given expression in radians. a. sin-(sin b. cos-(cos
Answer #1

5- (a) Here given, sin (sinh ) = siño (13) 2 3 [cos ( 2 ) = 773] [:sinh [ sin ( 3 ) = 27/33 Therefore required, siñ (sin 7• All-sin-tan-cos rule: If multipryer of It is eren, there is (tre) QI no change. If it is sin and cosec (tre) QII tan and co